PLZ SOLVE THIS
$8829.60 was the amount required to pay off a loan after 18 months.If the loan was at 8.80% per annum simple interest, how much of the total was interest? (Do not round your intermediate calculations and round the final answer to the nearest cent.)

Found 2 solutions by FrankM, addingup:
Answer by FrankM(1040)

8.8% simple interest per year is 13.2% over the 18 months period.
$8829.60/1.132 = $7800 to start.

Answer by addingup(3677)

8829.60= x(1+((0.088/12)*18))
8829.60= x(1+0.132)
8829.60= x1.132
x= 8829.60/1.132= 7,800 this is the original amount of the loan, and the interest:
8,829.60-7,800= 1,029.60 is the total interest.